OpenStudy (anonymous):

Use trigonometric identities to simplify the trigonometric expression (cscx+cotx)(cscx-cotx)

terenzreignz (terenzreignz):

Well first, we recall that \[\large (a+b)(a-b) = a^2 - b^2\] Analogous to this, what would be your first step in simplifying?

OpenStudy (anonymous):

umm would I start by dividing

terenzreignz (terenzreignz):

No :) Just recall this: \[\large (a+b)(a-b) = a^2 - b^2\] The product of the sum and difference of two numbers is the difference of their squares... It's like your a = csc(x) and b = cot(x) Now, do the first step in simplifying...

OpenStudy (anonymous):

canceling out right?

terenzreignz (terenzreignz):

What cancels out?

OpenStudy (anonymous):

(a-b)

OpenStudy (anonymous):

maybe?

OpenStudy (ybarrap):

use the hint @terenzreignz provided "a = csc(x) and b = cot(x)"

OpenStudy (dan815):